OpenStudy (anonymous):

If an 10 kg object has a potential energy of 125 J, what velocity would it be right before it hit the ground? Answer 250 m/s 50 m/s 25 m/s 5 m/s

OpenStudy (anonymous):

Hi when object hit to ground all of it's energy is KE & from conservation of energy we have PE(at first)=KE(in the end)so we have:\[mgh=(1/2)mv^2=125\rightarrow v=5\]

OpenStudy (anonymous):

see this:\[0.5\times10 \times v^2=125\rightarrow 10\times v^2=250\rightarrow v^2=25\rightarrow v=\sqrt{25}\]so v=5

OpenStudy (anonymous):

Wow. Yes. You are correct. Brain fart.